OpenStudy (anonymous):

The formula A = P + Prt represents the value, A, of an investment of P dollars at a yearly simple interest rate, r, for t years. The equation to model the value, A, of an investment of $54 at 9.26% for t years is given by A = 54 + 5t. The equation to model the value, A, of an investment of $84 at 2.38% for t years is given by A = 84 + 2t. Assuming A has the same value, the given equations form a system of two linear equations. Solve this system using an algebraic approach and interpret your answer. a.t = 5 The

OpenStudy (anonymous):

The formula A = P + Prt represents the value, A, of an investment of P dollars at a yearly simple interest rate, r, for t years. The equation to model the value, A, of an investment of $54 at 9.26% for t years is given by A = 54 + 5t. The equation to model the value, A, of an investment of $84 at 2.38% for t years is given by A = 84 + 2t. Assuming A has the same value, the given equations form a system of two linear equations. Solve this system using an algebraic approach and interpret your answer. a. t = 5 The two investments will reach the same value in 5 years. c. t = 1000 The two investments will reach the same value in 1000 years. b. t = 20 The two investments will reach the same value in 20 years. d. t = 10 The two investments will reach the same value in 10 years.

OpenStudy (anonymous):

@lexikaylynn @kidrah69 @Quan99 @autumnm @brittanydosey @chycora @damonte212 @Elephant16

OpenStudy (anonymous):

@butterflydreamer @RebeccaPink6398 @Chiko_1278 @basilalter

OpenStudy (damonte212):

Subtract the first equation from the second. The A's cancel out.

OpenStudy (damonte212):

3t = 84-54.

OpenStudy (damonte212):

ann thats your answer

OpenStudy (anonymous):

so it will be 10?